OpenStudy (anonymous):

Here's what I have done: (5,3)x(7,-2)/|(7,-2)| = 29/sqrt53 = 3.98 I thought I was done, but it says there is more to the problem. What do I do from here???

OpenStudy (amistre64):

\[proj_a b=\frac{\vec a.\vec b}{|a|^2}\ \vec a\]

OpenStudy (shayaan_mustafa):

@rebbylack look carefully. and understand what amistre64 says. He is great in mathematics. good luck.

OpenStudy (amistre64):

(5,3) (7,-2) : |..| = sqrt(53) -------- 35-6 = 29 \[\frac{29}{53}<7,-2>\]

OpenStudy (amistre64):

in effect, we are taking a unit vector in the direction of a and scaling it by the value |b| cos(theta)

OpenStudy (amistre64):

this gives us then the length and the direction we want 

OpenStudy (amistre64):

\[\cos\theta =\frac{a.b}{|a||b|}\] \[|b|\cos\theta =|b|\frac{a.b}{|a||b|}=\frac{a.b}{|a|}\] scale this in the direction of the unit vector of a \[|b|\cos\theta \frac{a}{|a|}=\frac{a.b}{|a|}\frac{a}{|a|}\to\ \frac{a.b}{|a|^2}\ a \]

OpenStudy (anonymous):

I think I'm beginning to understand... but I found the solution in my book and it says to use \[\frac{3.98}{\sqrt{53}} <5,3>\] ??? Does this way and your way both work?

OpenStudy (anonymous):

Oh wait...I see it now!!! they both equal the same thing. amistre64, you're a genius!! thank you sooo much :)

OpenStudy (amistre64):

you stated you want 5,3 ONTO 7,-2 so it important that your direction of the projection is in line with 7,-2 otherwise its not a projection ONTO 7,-2

OpenStudy (amistre64):

your last post has this thing running with <5,3> instead of <7,-2>